# On Multilingual Encoder Language Model Compression for Low-Resource Languages

### v1

@@ 6 Conclusion
We present an effective compression pipeline for multilingual encoder models designed for low-resource languages. By integrating staged knowledge distillation, structured pruning, hidden size truncation, and vocabulary reduction, we compress models by up to 92% while maintaining competitive performance - typically within 2–10% of the original - on four downstream tasks.
Limitations and Future Work
Future research should explore optimal strategies for transferring intermediate hidden size knowledge, possibly by applying structured pruning before distillation and identifying the most important components to keep instead of simply truncating them. Additionally, identifying language-specific layers and neurons and leveraging them for more targeted pruning could enhance compression efficiency while maintaining performance. Lastly, more advanced knowledge distillation methods should be experimented with.
Acknowledgements
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005).
@@ Appendix B Initialization Strategies for Knowledge Distillation
Figure 7: Validation accuracy comparing SVD and first-$k$ truncation initializations for mBERT models with reduced hidden size ($k$=312). Results show performance with two distillation approaches: using a layer-compressed model (Small teacher) or the original language-adapted model (Big teacher). Student models were trained with identical initialization methods but without knowledge distillation.
@@ Appendix F Knowledge Distillation Data Sizes
Table 2: Dataset sizes for knowledge distillation (KD) and monolingual fine-tuning (FT) for each language. The language-adapted models are sourced from Gurgurov et al. (2025), and the FT data sizes are as reported by them.
@@ Appendix I Adapter Trainable Parameter Counts
To examine whether the constrained task adapter capacity, as shown in Table 5, impacts downstream performance in compressed models, we vary the reduction factor $r$, thereby increasing adapter size (see Figure 10). We train task adapters on top of both full adapted models and hidden-size reduced models (564, 456, and 312). For the smallest models (456 and 312), we observe that increasing adapter capacity ($r$=2) leads to improved performance. However, this increase is unnecessary for larger mBERT variants (full and 564), while still beneficial for all small XLM-R models. These results suggest that for smaller models, increasing adapter capacity can yield modest performance gains. Tables 1 and 6 report results using the default reduction rate of 16.
@@
Model Configuration Reduction Size mBERT Avg.
mt sk sw
TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑ TC↑ SA↑ NER↑ POS↑
Base 0% 179M 68.73 65.82 59.96 89.04 85.31 91.98 91.44 97.01 69.60 64.62 83.81 87.61 79.58
Base-[mt, sk, sw] 0% 179M 84.91 73.55 65.00 93.96 86.26 91.87 90.42 96.87 86.68 81.28 82.49 88.69 85.17
KD layer red. ×2 -25% 135M 80.14 73.91 58.96 93.17 85.42 90.38 87.42 96.87 82.78 77.27 80.66 88.42 82.95
inter. layer red. → 2048 -30% 126M 79.03 74.74 58.13 92.74 85.25 90.23 88.48 96.67 83.21 75.90 79.76 88.54 82.72
* KD hid. size red. → 564 -50% 90M 79.47 70.22 61.12 92.64 83.36 90.48 88.07 96.30 83.50 76.10 79.74 88.42 82.45
vocab. red. → 40k -75% 45M 80.17 70.76 61.07 92.46 83.50 90.72 87.73 96.26 84.34 76.01 80.26 88.63 82.66
* KD hid. size red. → 456 -60% 71M 80.23 70.05 57.18 92.13 83.94 90.40 87.49 95.86 85.08 78.57 80.28 88.25 82.45
vocab. red. → 40k -80% 35M 80.95 69.72 55.87 92.00 84.22 90.39 87.38 95.95 83.03 78.50 79.79 88.37 82.18
* KD hid. size red. → 312 -73% 48M 73.14 71.98 39.50 90.33 80.91 90.40 86.48 95.48 81.75 76.48 79.56 87.72 79.48
vocab. red. → 40k -87% 23M 72.98 72.06 40.41 90.18 81.85 90.07 86.19 95.32 81.65 75.97 77.06 87.66 79.28
Table 6: Performance comparison before and after model compression for mBERT across Maltese (mt), Slovak (sk), and Swahili (sw). Task metrics: TC = Text Classification, SA = Sentiment Analysis, NER = Named Entity Recognition, POS = Part-of-Speech Tagging. All scores are averaged over three independent runs with different random seeds.

### v2

@@ 6 Conclusion
We present an effective compression pipeline for multilingual encoder models designed for low-resource languages. By integrating staged knowledge distillation, structured pruning, hidden size truncation, and vocabulary reduction, we compress models by up to 92% while maintaining competitive performance, typically within 2–10% of the original for moderate compression and 8–13% at maximum compression, on four downstream tasks.
Limitations
Our evaluation is limited to three low-resource languages and four downstream tasks, which may affect generalizability to other languages and task types. The compression pipeline requires target-language data for teacher adaptation, making it less suitable for truly low-resource languages with minimal corpora. We focus exclusively on encoder-only models (mBERT and XLM-R), and our structured pruning only targets feed-forward layers, leaving attention head pruning unexplored due to performance degradation.
Acknowledgments
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005), and by lorAI - Low Resource Artificial Intelligence, a project funded by the European Union under GA No.101136646.
@@ Appendix B Initialization Strategies for Knowledge Distillation
Figure 7: Validation accuracy comparing SVD vs. first-$k$ truncation for hidden size reduction to 312. “Small teacher” refers to the layer-compressed (6-layer) model; “Big teacher” is the original 12-layer language-adapted model. Truncation consistently outperforms SVD regardless of teacher size.
@@ Appendix F Knowledge Distillation Data Sizes
Table 2: Dataset sizes for knowledge distillation (KD) and monolingual fine-tuning (FT) for each language. The language-adapted models are sourced from gurgurov2025smallmodelsbigimpact, and the FT data sizes are as reported by them.
@@ Appendix I Adapter Trainable Parameter Counts
To examine whether the constrained task adapter capacity, as shown in Table 5, impacts downstream performance in compressed models, we vary the reduction factor $r$, thereby increasing adapter size (see Figure 10). We train task adapters on top of both full adapted models and hidden-size reduced models (564, 456, and 312). For the smallest models (456 and 312), we observe that increasing adapter capacity (rr=2) leads to improved performance. However, this increase is unnecessary for larger mBERT variants (full and 564), while still beneficial for all small XLM-R models. These results suggest that for smaller models, increasing adapter capacity can yield modest performance gains. Tables 1 and 6 report results using the default reduction rate of 16.
@@
Compression Stage Params Size Task Performance (F1) Avg
Maltese Slovak Swahili
TC SA NER POS TC SA NER POS TC SA NER POS
Baselines
   Multilingual 179M 179M 68.7 65.8 60.0 89.0 85.3 92.0 91.4 97.0 69.6 64.6 83.8 87.6 79.6
[…]
Compression Pipeline (minimal degradation)
   Layer reduction 135M (-25%) 135M 80.1 73.9 59.0 93.2 85.4 90.4 87.4 96.9 82.8 77.3 80.7 88.4 83.0
   + FFN pruning 126M (-30%) 126M 79.0 74.7 58.1 92.7 85.3 90.2 88.5 96.7 83.2 75.9 79.8 88.5 82.7
   + Hidden 564 90M (-50%) 90M 79.5 70.2 61.1 92.6 83.4 90.5 88.1 96.3 83.5 76.1 79.7 88.4 82.5
[…]
Further compression (moderate degradation)
   + Hidden 456 71M (-60%) 71M 80.2 70.1 57.2 92.1 83.9 90.4 87.5 95.9 85.1 78.6 80.3 88.3 82.5
   + Vocabulary 35M (-80%) 35M 81.0 69.7 55.9 92.0 84.2 90.4 87.4 96.0 83.0 78.5 79.8 88.4 82.2
Maximum compression (higher degradation)
[…]
Table 6: Progressive compression of mBERT. Stages are grouped by degradation level. Highlighted rows indicate the baseline (gray) and optimal compression point (green, 75% reduction with 2.5% drop). Maximum compression rows (red) show significant degradation (5.9% drop). TC=Topic Classification, SA=Sentiment Analysis, NER=Named Entity Recognition, POS=Part-of-Speech Tagging. F1 scores averaged over 3 runs.